OpenStudy (anonymous):

Find the number of permutations of the first 8 letters of the alphabet taking 5 letters at a time. A. 1,680 B. 56 C. 6,720 D. 109

OpenStudy (shamim):

\[P _{5}^{8}=\frac{ 8! }{ (8-5)! }\]

OpenStudy (shamim):

did u get it

OpenStudy (shamim):

u know\[2!=2 \times 1\]

OpenStudy (shamim):

\[3!=3 \times 2 \times 1\]

OpenStudy (shamim):

so\[8!=8 \times 7 \times 6 \times 5 \times 4 \times 3 \times 2 \times 1\]